OpenStudy (anonymous):

Write an equation that passes through the points (4,-1) and (2,- 2).

OpenStudy (raden):

use the formula : (y-y1)/(y2-y1) = (x-x1)/(x2-x1)

OpenStudy (anonymous):

-1/-2

OpenStudy (anonymous):

i did, y2-y1/x2-x1

OpenStudy (raden):

yes, that's just for the slope not whole like ur question ask

OpenStudy (anonymous):

lol i dont understand the y-y1 part .... does that mean y-(-1)/-2-(-1) ?

OpenStudy (johnweldon1993):

@RadEn That equation above is new to me! O.o but awesome! thanks for posting that!

OpenStudy (raden):

well, another way is use the formula of a straight line : y = mx + c (with m is the slope) you got m = 1/2, put m = 1/2 into equation above : y = 1/2 x + c then take one of the points given above, let it is point (4,-1) subtitute that point then solve for c y = 1/2 x + c -1 = 1/2 * 4 + c -1 = 2 + c c = -3 finally, we have an equation completely it is y = 1/2 x - 3

OpenStudy (anonymous):

WOAH! thats a brain scrambler hahah, thanks for helping me out. You might need to stick around more. :D

OpenStudy (raden):

really, i got it in my book since 20 year ago, lol @johnweldon1993

OpenStudy (johnweldon1993):

haha, no the way I learned was to find the slope (y2-y1)/(x2-x1) then plug that into the point slope formula.... now, mind = blown! haha thanks for that :D wonder what else I've missed lol

OpenStudy (raden):

hehe, thanks :)

OpenStudy (johnweldon1993):

Now I'm going to go through my calc books again and make sure I'm not overlooking something that would make life 20X easier! haha

OpenStudy (raden):

:)